# Real-time Free-view Human Rendering from Sparse-view RGB Videos using Double Unprojected Textures

### v1

@@ 5 Discussion
Limitations. Although the proposed method, DUT, achieves unprecedented rendering quality, speed, and robustness for sparse human capture and free-view rendering, there are remaining challenges opening up directions for future work. First, our method cannot handle extremely loose cloth, like skirts. Here, recent advances in cloth simulation  [13, 28] combined with our rendering paradigm might be a promising solution. Second, our model relies on reasonable motion capture results as inputs. Here, deeply entangling the motion capture component with the rendering task can offer a potential solution. Third, the deformation model relies on scans or point-clouds for supervision, which require time-consuming per-frame reconstruction. Involving silhouette supervision [20] or foundational depth priors [85] may alleviate the need for this step.
@@ Appendix A Overview
We present more details and experimental results of DUT towards a more clear understanding and in-depth analysis. We first show an interactive demo with our method (Sec. B). Then, we offer additional details about the visibility computation (Sec. C), loss functions (Sec. D), implementation details (Sec. E), and runtime analysis (Sec. F). We also provide additional experimental results about robustness with respect to motion capture qualities (Sec. G) and out-of-distribution motions (Sec. H). Tab. 7 lists notations and symbols used in the main paper, and demonstrates their descriptions.
@@ Appendix D Loss Function
3090 1 K 256 3.8723 7.6027 11.4430 14.6424 21.3143 23.5674
3090 4 K 256 3.8751 7.6097 11.5076 14.7454 21.4354 23.7146
3090 1 K 512 3.8669 7.5927 16.9868 20.3320 27.0508 29.4855
3090 4 K 512 3.8601 7.5768 16.9993 20.3178 27.0191 29.3762
H100 1 K 256 3.8130 6.9313 10.7382 13.3484 16.5124 18.7570
H100 4 K 256 3.8489 6.9685 10.7392 13.3512 16.4705 18.7303
H100 1 K 512 3.7821 6.8754 12.9675 15.6070 18.7620 21.0133
H100 4 K 512 3.8300 6.8982 12.9877 15.6178 18.7669 21.0022
Table 4: Quantitative Ablation. Here, we demonstrate a detailed runtime ablation of our method with different texture resolutions (Tex Res), rendering resolutions (Rnd Res), and GPUs. All timings in the table are the accumulated timings and the unit is milliseconds.
@@ Appendix E Implementation Details
Training & Testing Actions

### v2

@@ 5 Discussion
Limitations. Although the proposed method achieves unprecedented rendering quality, speed, and robustness for sparse human capture and free-view rendering, there are remaining challenges opening up directions for future work. First, our method can not maintain the rigidity of the fingers. We can improve finger geometry by removing deformations on hands [59] or estimating global hand translations. Second, our method inherits the limitation of template-based methods [19, 52, 59] in that it is hard to handle topological changes, like unzipping a jacket. Third, the deformation model relies on scans or point-clouds for supervision, which require time-consuming per-frame reconstruction. Involving silhouette supervision [19] or foundational depth priors [83] may alleviate the need for this step.
@@
Acknowledgements This project was supported by the Saarbrücken Research Center for Visual Computing, Interaction and AI, as well as the DFG’s Research Training Group on “Neuroexplicit Models of Language, Vision, and Action” (GRK 2853/1).
@@ Appendix A Overview
We present more details and experimental results of DUT towards a more clear understanding and in-depth analysis. We first show an interactive demo with our method (Sec. B). Then, we offer additional details about the visibility computation (Sec. C), loss function (Sec. D), implementation details (Sec. E), runtime analysis (Sec. F), undeformed texture maps (Sec. L) and limitation discussions (Sec. M). We also provide additional experimental results about influence with respect to motion capture qualities (Sec. G), out-of-distribution motions (Sec. H), motion sensitivity analysis (Sec. I), performance on a loose and long hair subject (Sec. J), novel lighting (Sec. K). Tab. 7 lists notations and symbols used in the main paper, and demonstrates their descriptions.
@@ Appendix D Loss Function
3090 1 K 256 3.8723 7.6027 11.4430 14.6424 21.3143 23.5674 42.43
3090 4 K 256 3.8751 7.6097 11.5076 14.7454 21.4354 23.7146 42.17
3090 1 K 512 3.8669 7.5927 16.9868 20.3320 27.0508 29.4855 33.91
3090 4 K 512 3.8601 7.5768 16.9993 20.3178 27.0191 29.3762 34.04
H100 1 K 256 3.8130 6.9313 10.7382 13.3484 16.5124 18.7570 53.31
H100 4 K 256 3.8489 6.9685 10.7392 13.3512 16.4705 18.7303 53.39
H100 1 K 512 3.7821 6.8754 12.9675 15.6070 18.7620 21.0133 47.59
H100 4 K 512 3.8300 6.8982 12.9877 15.6178 18.7669 21.0022 47.61
Table 4: Quantitative Ablation. Here, we demonstrate detailed runtime ablation of our methods with different texture resolutions (Tex Res), rendering resolutions (Rnd Res) and GPUs. All the time in the table is the accumulated time from the beginning and their units are milliseconds.
@@ Appendix E Implementation Details
Training & Testing Action
@@ Appendix F Runtime Analysis
Figure 12: Motion Sensitivity Analysis to Increasing Errors. Our method still outputs reasonable results to different level of motion capture errors at inference.
Figure 13: Qualitative results. We perform our method on a loose and long-hair subject, it manages to capture the coarse deformations of hair and dress and produces faithful rendering results.
Figure 14: Qualitative results. Results with finetuned model under novel lighting. After finetuning, DUT still runs in feed-forward manner.
Figure 15: Illustration of Undeformed Texture Map. The distortions of undeformed (first) texture maps are directly related with deformations on the canonical template.
Figure 16: Topology Change. Results of taking off cloth.
@@ Appendix H Additional Comparisons on OOD Motions
Appendix I Sensitivity Analysis to Motion Errors
We show a sensitivity analysis by manually adding linear motion capture errors to left elbow from 0 to $60^{\circ}13^{{}^{\prime}}$ in Fig. 12. Surprisingly, DUT produces good results within $7^{\circ}46^{{}^{\prime}}$ error (0, $3^{\circ}43^{{}^{\prime}}$, $7^{\circ}46^{{}^{\prime}}$) and reasonable results even with $60^{\circ}13^{{}^{\prime}}$ error.
Appendix J Performance on Loose Clothing and Long Hair Subject
In Fig. 13, we show a subject with dynamic long-hair wearing a dress. Our geometry module recovers the coarse geometry of hair dynamics and dress. After which, our rendering module faithfully generates high fidelity renderings.
Appendix K Results under Novel Lighting Environment
Towards realistic application scenarios, once our method is trained, we may run it with novel cameras under novel lighting environments. We can either perform color augmentation to the input views during training [79] or finetune the model on the sparse input views captured under the new illumination. We provide a preliminary result for this in Fig. 14 where we finetuned the model on the input views and isolated frames (not used in testing) in a new lighting condition. After finetuning, DUT still runs in feed-forward manner. Our method can be easily adopted to the novel lighting and produce reasonable rendering results.
Appendix L Additional Discussions about Undeformed (First) Texture Map
Our GeoNet $\Phi_{\mathrm{Geo}}$ estimates template deformations from undeformed texture map $\mathcal{T}_{\mathrm{c,1st}}$ and non-root normal map $\mathcal{T}_{\mathrm{\bar{N}}}$. Similar to the normalized skeletal motion in DDC [19], $\mathcal{T}_{\mathrm{\bar{N}}}$ provides the pose information. However, dynamic geometries of a moving human body are not completely determined by the skeletal pose at that moment, it leads to one-to-many mapping issue [40]. Our undeformed texture map $\mathcal{T}_{\mathrm{c,1st}}$ offers additional information about the degree of deformations. As shown in Fig. 15, under the same body pose, the degree of deformations can be reflected by the distortions of $\mathcal{T}_{\mathrm{c,1st}}$.
Appendix M More Discussions about Limitations
Color Fluctuation. Though simple and efficient, our method suffers from a certain extent color fluctuation of some frames as shown in the video. We found that this could be attributed to the predicted Gaussians are trying to overfit the uneven lightings of studio and shadows on the body, which are challenging for such simple representation. Integrating ray tracing or physically based rendering may reduce the color fluctuation.
Topology Change. In Fig. 16, we show results on a sequence where the subject is taking clothes off. While results look reasonable, still the quality degrades. Though a fixed template contributes a lot, it will be an interesting direction to investigate how to introduce dynamic template update into such task, especially with only RGB inputs.